### v1

Abstract.
Formulas of Harer-Zagier and Harder imply that the orbifold Euler characteristic of the moduli stacks of smooth curves or principally polarized abelian varieties have a sign given by $(-1)^{\dim{\mathscr{M}}_{g}}$ or $(-1)^{\dim{\mathscr{A}}_{g}}$. This is generalized as follows: Given a perverse sheaf on a product of moduli stacks of either type, it is proved that the Euler characteristic is nonnegative when the base field has characteristic zero. This is shown to be false in positive characteristic.
First author supported by a grant from the Simons Foundation
1. Introduction
Let ${\mathscr{M}}_{g,n}$ denote the moduli stack of smooth projective curves of genus $g$ and $n$-marked points, let ${\mathscr{A}}_{g}$ denote the moduli stack of principally polarized $g$-dimensional abelian varieties, and let ${\mathscr{A}}_{g,n}$ the $n$-fold fibre product of the universal family over ${\mathscr{A}}_{g}$. We work over an algebraically closed base field $k$. In this setting (and over $k={\mathbb{C}}$), we have the following results due to Harer-Zagier [15] and Harder [14]:
  1. (1)
    The orbifold characteristic or equivalently the Euler-Satake characteristic
    $\chi^{orb}({\mathscr{M}}_{g,n})=(-1)^{n}\frac{(2g-1)B_{2g}}{(2g)!}(2g+n-3)!\quad\text{when }g\geq 2$
  2. (2)
    If $n\geq 3$, then $\chi^{orb}({\mathscr{M}}_{0,n})=(-1)^{n-3}(n-3)!$.
  3. (3)
    If $n\geq 1$, then $\chi^{orb}({\mathscr{M}}_{1,n})=(-1)^{n}\frac{(n-1)!}{12}$
  4. (4)
    The orbifold Euler characteristic
    $\chi^{orb}({\mathscr{A}}_{g})=\zeta(-1)\zeta(-3)\cdots\zeta(1-2g).$
Corollary 1.1.
The sign of $\chi^{orb}({\mathscr{M}}_{g,n})$ is given by $(-1)^{\dim{\mathscr{M}}_{g,n}}$ and the sign of $\chi^{orb}({\mathscr{A}}_{g})$ is given by $(-1)^{\dim({\mathscr{A}}_{g})}$
Proof.
This follows from the standard formulas $\dim({\mathscr{M}}_{g,n})=3g-3+n$, $\dim({\mathscr{A}}_{g})=\frac{g(g+1)}{2}$, $\zeta(1-2m)=-\frac{B_{2m}}{2m}$, and
$\operatorname{sign}B_{2m}=\begin{cases}+1&\text{if $m$ odd}\\ -1&\text{if $m$ even}\end{cases}$
(We take $\operatorname{sign}x$ to be $1$ if $x\geq 0$ and $-1$ otherwise.) ∎
In this article, we generalize this corollary by showing that the sign of the Euler characteristic of any smooth closed irreducible substack ${\mathscr{X}}$ of one the previous stacks satisfies the same condition. In fact, we can strengthen this as follows. Given a Deligne-Mumford stack defined over $k$, let $\operatorname{Perv}({\mathscr{X}})$ the category of perverse sheaves. In the following, this will mean one of two things. If $k={\mathbb{C}}$, one can consider the underlying analytic Deligne-Mumford stack and perverse sheaves with ${\mathbb{Q}}$ coefficients (in the analytic topology). For an arbitrary arbitrary field $k$ with $char(k)=p$, one can consider the category of $\ell$-adic perverse sheaves (for some $\ell\neq p$). In this setting and when $char(k)=0$, one can define the Euler-Satake characteristic of a perverse sheaf (as a rational number). The precise definition of $\operatorname{Perv}({\mathscr{X}})$ and the notion of Euler characteristic for an object of this category is given in section 3. Let us start by stating a special case of our main results.
Theorem 1.2.
Let ${\mathscr{X}}$ be one of ${\mathscr{M}}_{g,n}$ (with $02g-2+n>$ 0) or ${\mathscr{A}}_{g,n}$, and suppose $char(k)=0$. Then for ${\mathcal{K}}\in\operatorname{Perv}({\mathscr{X}})$
$\chi^{orb}({\mathscr{X}},{\mathcal{K}})\geq 0.$
Corollary 1.3.
If ${\mathscr{X}}\subset{\mathscr{A}}_{g,n}$ or ${\mathscr{X}}\subset{\mathscr{M}}_{g,n}$ is a smooth closed irreducible substack, then
$(-1)^{\dim X}\chi^{orb}({\mathscr{X}})\geq 0.$
More generally, the analogous assertion holds for any local complete intersection.
Proof.
Take ${\mathcal{K}}=\iota_{*}{\mathbb{Q}}_{{\mathscr{X}}}[\dim{\mathscr{X}}]$ where $\iota$ is the inclusion of ${\mathscr{X}}$. This is a perverse sheaf by [18, chap III, lemma 6.5]. ∎
Corollary 1.4.
If ${\mathcal{K}}$ is a perverse sheaf on a Shimura variety ${\mathscr{X}}$ of Hodge type and finite level, then $\chi^{orb}({\mathscr{X}},{\mathcal{K}})\geq 0$.
Proof.
Essentially by definition, ${\mathscr{X}}$ will embed as closed smooth substack of ${\mathscr{A}}_{g}$ for some $g$. ∎
In particular, this theorem gives a simple proof recovering the sign of the Euler characteristic of ${\mathscr{M}}_{g,n}$ and ${\mathscr{A}}_{g,n}$. We expect that the last corollary should hold for arbitrary Shimura varieties.
Before we explain the basic idea of the proofs, we should note that the above theorem is in fact a combination of special cases of two separate theorems. The first result involves abelian varieties.
Theorem 1.5.
Let ${\mathcal{K}}$ be a perverse sheaf on ${\mathscr{A}}_{g_{1},n_{1}}\times\ldots\times{\mathscr{A}}_{g_{r},n_{r}}$, and suppose $char(k)=0$. Then
$\chi^{orb}(X,{\mathcal{K}})\geq 0.$
By using the moduli spaces $\prescript{}{N}{{\mathscr{A}}_{g,n}}$ with level $N\geq 3$ structures, we reduce to an analogous statement about varieties rather than stacks. To check this version of the theorem, we verify that the log cotangent bundle of a product of toroidal compactifications
$\overline{X}=\overline{{}_{N_{1}}{\mathscr{A}}_{g_{1},n_{1}}}\times\ldots\times\overline{{}_{N_{r}}{\mathscr{A}}_{g_{r},n_{r}}}$
is nef. This is reduced to a formula of Faltings and Chai, which expresses the log cotangent bundle of each factor as the second symmetric power of a Hodge bundle, and the Fujita-Kawamata semipositivity theorem for the latter. The above theorem now follows from the nefness result together with the log Dubson-Kashiwara theorem and an inequality of Fulton-Lazarsfeld. This last step of the proof is similar to, and inspired by, arguments used in recent work on the Hopf-Singer and related conjectures (see [4, 23]). We also note that the idea of using the Dubson-Kashiwara formula to show positivity of Euler charactersitic of perverse sheaves first appears (in the context of semi-abelian varieties) in the work of Franecki-Kapranov ([11]).
Our second result concerns moduli of curves. It is convenient to work with the moduli stack ${\mathscr{M}}_{g,n}^{c}\subset\overline{{\mathscr{M}}_{g,n}}$ of genus $g$ stable curves of compact type with $n$ marked points. Compact type means that the dual graph is a tree. In particular, ${\mathscr{M}}_{g,n}\subset{\mathscr{M}}_{g,n}^{c}$. Our second main theorem is the following:
Theorem 1.6.
Let ${\mathscr{X}}={\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{c}$ where $g_{i}>1$ or $g_{i}=1,n_{i}>0$, and suppose $char(k)=0$. If ${\mathcal{K}}$ is a perverse sheaf on ${\mathscr{X}}$, then
$\chi^{orb}({\mathscr{X}},{\mathcal{K}})\geq 0$
We will define a proper morphism
$\tau:{}_{N}{\mathscr{M}}_{g,n}^{c}\to{}_{N}{\mathscr{A}}_{g,n}$
which agrees with the Torelli map on ${}_{N}{\mathscr{M}}_{g}$. It easy to see that $\tau$ is a locally closed embedding of ${}_{N}{\mathscr{M}}_{g,n}$, but it is not an embedding of the whole space. So Theorem 1.6 does not follow immediately from Theorem 1.5. On the other hand, Theorem 1.5 does apply to perverse sheaves supported on the image $Y=\tau({}_{N}{\mathscr{M}}_{g,n}^{c})$. Now setting $X={}_{N}{\mathscr{M}}_{g,n}^{c}$ and $U={}_{N}{\mathscr{M}}_{g,n}$, we conclude the proof with the help of the following technical result proved in section 6 using Beilinson’s gluing construction (for perverse sheaves) and a deformation to the normal cone technique.
Theorem 1.7.
Suppose we are given a diagram of quasiprojective varieties
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tilde{j}}$$\scriptstyle{j}$${Y}$
where $j,\tilde{j}$ are open immersions, and $f$ is projective and birational, and $Z=X-U$ is a Cartier divisor. Suppose that for all $K^{\prime}\in\operatorname{Perv}(Y)$ and $K^{\prime\prime}\in\operatorname{Perv}(Z)$ we have $\chi(Y,K^{\prime})\geq 0$ and $\chi(Z,K^{\prime\prime})\geq 0$. Then for all $K\in\operatorname{Perv}(X)$, we have $\chi(X,K)\geq 0$.
In the final section of this paper, we show that the analog of Theorem 1.2 fails in positive characteristic. Let us now view $\prescript{}{N}{{\mathscr{A}}}_{g}$ as a scheme over $\operatorname{Spec}{\mathbb{Z}}[1/N]$, and let $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes\bar{{\mathbb{F}}}_{p}$ denote the fibre over $\bar{{\mathbb{F}}}_{p}$. We first observe that the ($\ell$-adic) Euler characteristic is indepenpent of $p\nmid N$ (see 9.5).
Theorem 1.8.
For all $p\nmid N$, $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes\bar{{\mathbb{F}}}_{p})=\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}}).$
On the other hand, the existence of supersingular strata allows one to construct counter-examples to our main theorem in characteristic p. In particular, we show in Corollary 9.6 that:
Theorem 1.9.
There exists a sequence of smooth subvarieties $V_{g_{i}}\subset\prescript{}{N}{{\mathscr{A}}}_{g_{i}}\otimes\bar{{\mathbb{F}}}_{p}$ such that $(-1)^{\dim V_{g_{i}}}\chi(V_{g_{i}})\to-\infty$.
This is closely related to the fact, first observed by Moret-Bailly [25], that the Fujita-Kawamata theorem fails in positive characteristic. See Remark 9.7. We note that we work with moduli with level structure, because it is not clear how to define the Euler-Satake characteristic of a Deligne-Mumford stack in characteristic $p$.
Finally, we briefly outline the contents of the various sections. The second section recall some background results which will be used in the following sections. In particular, we recall a positivity result of Fulton-Lazarsfeld, a log version of the Dubson-Kashiwara formula, and apply this to prove semi-positivity of certain Hodge bundles. We also recall a fundamental result of Beilinson on gluing perverse sheaves. In section 3, we give the definition of the Euler-Satake characteristic of construcitble sheaves on Deligne-Mumford stacks (following [5]), and recall some facts on the perverse sheaves on stacks. In particular, we observe that the constant sheaf on a local complete intersection stack is a perverse sheaf. In section 4, we prove the signed Euler characteristic property for ${}_{N}{\mathcal{A}}_{g}$ and their products. In section 5, we prove analogous assertions for the universal family over ${}_{N}{\mathcal{A}}_{g}$, and apply this to prove Theorem 1.5. In section 6, we prove Theorem 1.7. In section 7, we construct an extension of the Torelli map to ${\mathscr{M}}^{c}_{g,n}$. Finally, these results are applied in section 8 to prove Theorem 1.6 and finish the proof of Theorem 1.2. In section 9, we discuss the aforementioned results in characteristic p.
Acknowledgements: The first author thanks Laurentiu Maxim and Botong Wang from introducing him to some of these ideas in connection with the Hopf-Singer conjecture.
2. Background
In this section, we recollect some standard results which will be useful in the following.
2.1. Fulton-Lazarsfeld for nef bundles
Let $X$ be a smooth projective variety, and $E$ be a vector bundle. A conical subset of $E$ is a Zariski closed subset invariant under the natural ${\mathbb{G}}_{m}$-action. In this setting, one has the following positivity of intersection products result due to Fulton-Lazarsfeld.
Theorem 2.1.
Suppose $E$ is nef of rank $e$. Then for any conical subset $C\subset E$ of dimension $e$,
$\langle C,X\rangle\geq 0$
where $\langle C,X\rangle$ is the degree of the zero cycle given by the intersection of $C$ with the zero section.
Proof.
See [21], 8.2.6. ∎
2.2. Log Dubson-Kasiwara
Let $X$ be a smooth complex variety of dimension $d$, and ${\rm D}^{b}_{c}(X)$ denote the bounded derived category of constructible sheaves (of ${\mathbb{C}}$-vector spaces) on $X$. Given ${\mathcal{K}}\in{\rm D}^{b}_{c}(X)$, there is a well-defined cycle $CC({\mathcal{K}})\in Z^{d}(T^{*}{X})$, where $Z^{d}(T^{*}(X)$ is the group of cycles of codimension $d$. This construction satisfies the following well known properties which we collect below:
  1. (1)
    The components of $CC({\mathcal{K}})$ are conic Lagrangian subsets. In particular, they are of the form $\overline{T^{*}_{Z_{reg}}X}\subset T^{*}{X}$ where $Z\subset X$ is a closed subvariety and $Z_{reg}$ denotes the regular locus.
  2. (2)
    If ${\mathcal{P}}$ is a perverse sheaf on $X$, the $CC({\mathcal{P}})$ is an effective cycle.
  3. (3)
    (Dubson-Kashiwara Formula) Suppose furthermore, that $X$ is projective. Then, one has
    $\chi(X,{\mathcal{K}})=\langle CC({\mathcal{K}}),X\rangle.$
    Here the right hand side is the degree of the zero cycle given by the intersection of the characteristic cycle and the zero section.
In ([29]), Wu and Zhou give an extension of the Dubson-Kashiwara formula to the setting of log cotangent bundles. Suppose now that $(X,D)$ is a pair where $X$ is a smooth projective variety and $D$ is a simple normal crossings divisor. In this case, one have the standard exact sequence:
$0\rightarrow\Omega^{1}_{X}\rightarrow\Omega^{1}_{X}(\log D)\rightarrow{\mathcal{O}}_{D}\rightarrow 0$
In particular, this gives an injective map of the corresponding vector bundles
$T^{*}X\hookrightarrow T^{*}X(\log D).$
If $U:=X\setminus D$, then we obtain an open immersion
$T^{*}U\hookrightarrow T^{*}X(\log D).$
Given ${\mathcal{K}}\in{\rm D}^{b}_{c}(U)$, let $\overline{CC({\mathcal{K}})}$ denote its closure in $T^{*}X(\log D)$. This is a conical subset of dimension $d$. In [29, theorem 1.6], the authors show that
(2.1.1) $\chi(U,{\mathcal{K}})=\langle\overline{CC({\mathcal{K}})},X\rangle.$
2.3. Kähler hyperbolicity
Given a Kähler manifold $X$, the universal cover $\pi:\tilde{X}\to X$ carries an induced Kähler metric. The manifold $X$ is Kähler hyperbolic if the pullback of the Kähler form $\pi^{*}\omega=d\alpha$, where $\alpha$ has bounded norm with respect to the induced metric. We have the following examples [13, ex 0.3A]:
  1. (1)
    A Hermitian locally symmetric space of noncompact type is Kähler hyperbolic.
  2. (2)
    A complex submanifold of a Kähler hyperbolic manifold is Kähler hyperbolic.
2.4. Log structures
It is convenient to formulate some results in the language of log structures. We will not define them here, but instead refer the reader to Kato’s paper [16] for details. Let $k$ denote a field of characteristic 00. Given a smooth $k$-variety $Y$, a divisor $D\subset Y$ with reduced normal crossings induces a log structure [16, ex 1.5], which we denote by $(Y,D)$. We refer to a log structure of this type as having normal crossing type or just nc. If $f:X\to Y$ is a morphism of smooth varieties, with reduced divisors with normal crossings $D\subset Y$ and $E\subset X$ such that $f^{-1}D\subseteq E$, then we get morphism of nc log schemes $f:(X,E)\to(Y,D)$. We say $f$ is log smooth if it is a smooth morphism of log schemes [16, §3.3]. This implies that
$\Omega^{1}_{(X,E)/(Y,D)}:=\Omega_{X}^{1}(\log E)/f^{*}\Omega_{Y}^{1}(\log D)$
is locally free of rank equal to the relative dimension of $X/Y$. We note that nc log schemes are log smooth over $\operatorname{Spec}(k)$ (where $\operatorname{Spec}(k)$ has the trivial log structure).
2.5. Hodge bundles
Suppose that $(Y,D)$ is a projective nc log scheme over $k$ as above. Let $\overline{f}:(X,E)\to(Y,D)$ be a projective log smooth morphism to $(Y,D)$, which restricts to an abelian scheme $f:A\to Y-D$ over $Y-D$. Let $\overline{F}^{1}=\overline{f}_{*}\Omega^{1}_{(X,E)/(Y,D)}$.
Theorem 2.2 (Fujita-Kawamata).
The sheaf $\overline{F}^{1}$ is a nef vector bundle.
Proof.
Without loss of generality, we can assume that $k={\mathbb{C}}$. Switching to the analytic topology, $R^{1}f^{an}_{*}{\mathbb{C}}$ underlies a polarized variation of Hodge structure $(V,F^{\bullet},\ldots)$ of type $\{(1,0),(0,1)\}$, with unipotent monodromy about components of $D$. The Hodge bundle $F^{1}V=f^{an}_{*}\Omega^{1}_{A/Y-D}$ extends to $(\overline{F}^{1})^{an}$ over $Y$. The theorem now follows from the much more general result due to Fujino-Fujisawa-Saito [12, theorem 3]. ∎
We will refer to $\overline{F}^{1}$ as the Hodge bundle associated to $f:A\to Y-D$.
2.6. Beilinson’s gluing construction
We recall a basic result of Beilinson [6, 24, 27]. We will follow the notation from [7] and write $f_{*}$ etc. for the derived direct image. Also, $\psi_{f},\phi_{f}$ denotes derived nearby/vanishing cycle functors shifted by $[-1]$ (so as to preserve perverse sheaves). Let $\psi^{u}_{f}\subset\psi_{f},\phi^{u}_{f}\subset\phi_{f}^{u}$ denote the maximal subfunctors on which the monodromy $T$ acts unipotently. The $can$ and $var$ maps restrict to these. Let $X$ be a variety with a regular function $g:X\to{\mathbb{A}}^{1}$. Let $Z:=g^{-1}(0)$, $U=X-Z$, and let $i:Z\hookrightarrow X,j:U\hookrightarrow X$ denote the canonical inclusions. In this setting, one has the following fundamental result due to Beilinson.
Theorem 2.3 (Beilinson).
There is an equivalence between the category of perverse sheaves $\operatorname{Perv}(X)$ and the category
$\begin{split}B_{g}(U,Z)&=\{(K_{U},K_{Z},c,v)\mid K_{U}\in\operatorname{Perv}(U),K_{Z}\in\operatorname{Perv}(Z)\\ &c:\phi_{g}^{u}K_{U}\to K_{Z},v:K_{Z}\to\phi_{g}^{u}K_{U},cv=T-I\}\end{split}$
In one direction, $K\in\operatorname{Perv}(X)$ goes to $(j^{*}K,\phi_{g}^{u}K,can,var)$. The inverse is more complicated. Beilinson constructs a “maximal extension” functor $\Xi_{g}:\operatorname{Perv}(U)\to\operatorname{Perv}(X)$ which fits into exact sequences
[…]
The inverse sends an object $(K_{U},K_{Z},c,v)$ to the cohomology of the complex
(2.3.2) $K=H^{0}(i_{*}\psi^{u}_{g}K_{U}\xrightarrow{(c,\alpha)}\Xi_{g}K_{U}\oplus i_{*}K_{Z}\xrightarrow{(v,-\beta)}i_{*}\psi^{u}_{g}K_{U})$
(where this is centered in degree 0).
Remark 2.4.
We observe for later that the first map in (2.3.2) is injective, and the second is surjective.
2.7. Monodromic Sheaves
Let $X$ be a variety over an algebraically closed field $k$ such that $char(k)=0$. Suppose that $X$ comes equipped with a ${\mathbb{G}}_{m}$-action denoted by $\theta:{\mathbb{G}}_{m}\times X\rightarrow X.$ For $\lambda\in k^{\times}$, let $\theta_{\lambda}:=\theta(-,\lambda):X\rightarrow X$, and for $x\in X(k)$, let $w_{x}=\theta(x,-):{\mathbb{G}}_{m}\rightarrow X$.
Example 2.5.
A basic example of such data arises from a cone. More precisely, given a variety $X$ as above and a quasi-coherent graded ${\mathcal{O}}_{X}$-algebra ${\mathcal{A}}=\oplus_{i\geq 0}{\mathcal{A}}_{i}$ of finite type generated in degree 1 such that ${\mathcal{A}}_{0}={\mathcal{O}}_{X}$, the usual $\operatorname{Spec}$ construction gives rise to an affine morphism $\operatorname{Spec}({\mathcal{A}})\rightarrow X$. The scheme $\operatorname{Spec}({\mathcal{A}})$ has a natural ${\mathbb{G}}_{m}$-action coming from the grading. Most of our examples will arise in this fashion.
A constructible sheaf (respectively complex of sheaves) ${\mathcal{K}}$ on $X$ is monodromic if $w_{x}^{*}({\mathcal{K}})$ (respectively the cohomology sheaves of $w_{x}^{*}({\mathcal{K}})$) are locally constant for all $x\in X(k)$. We let $Perv_{m}(X)$ denote the full-subcategory of monodromic perverse sheaves.
Lemma 2.6.
([28], 3.2) With notation as above, the following are equivalent:
  • For all $\lambda\in k^{\times}$, $\theta_{\lambda}^{*}{\mathcal{K}}\cong{\mathcal{K}}$.
  • ${\mathcal{K}}$ is monodromic.
Proof.
In the setting of cones and constructible sheaves, this is contained Proposition 3.2 of [28]. The same proof goes through in the current setting. Moreover, the case of complexes can be reduced to sheaves by passing to cohomology sheaves. ∎
Let $X$ be a variety equipped with a ${\mathbb{G}}_{m}$-action, and $h:X\rightarrow{\mathbb{A}}^{1}$ a morphism which is equivariant for the standard ${\mathbb{G}}_{m}$-action on ${\mathbb{A}}^{1}$. Then $Z:=h^{-1}(0)$ and $U$ are invariant under ${\mathbb{G}}_{m}$. In this case, we may consider the categories $\operatorname{Perv}_{m}(X),\operatorname{Perv}_{m}(U),$ and $\operatorname{Perv}_{m}(Z)$.
Theorem 2.7.
With notation as above, $\operatorname{Perv}_{m}(X)$ is equivalent to the full sub-category $B_{g,m}(U,Z)\subset B_{g}(U,Z)$ consisting of objects such that $K_{U}$ and $K_{Z}$ are monodromic.
Proof.
First, note that $j^{*}{\mathcal{K}}$ is monodromic if ${\mathcal{K}}$ is monodromic. This follows directly from the previous lemma. Moreover, the vanishing cycles along $h$ of a monodromic sheaf is also monodromic. This follows from the previous lemma and the fact that vanishing cycles commutes with pull-back along acyclic morphisms. Since smooth morphisms and isomorphisms are such morphisms the result follows from the previous lemma (apply the previous remarks to $\theta_{\lambda}$). ∎
Remark 2.8.
Suppose that $f:X\rightarrow Y$ is compatible with ${\mathbb{G}}_{m}$-actions. Then $f_{!}$ preserves monodromic sheaves. This follows from the previous lemma and the fact that $f_{!}$ is compatible with base change. Namely, one has the cartesian square:
[…]
In particular, this holds for $f_{*}$ when $f$ is proper, and $j_{!}$ where $j$ is an open immersion.
3. Perverse sheaves on Stacks and Euler characteristics
In this section, we explain the construction of the Euler-Satake charactersitic for constructible sheaves on Deligne-Mumford or DM stacks, and recall some basic facts about perverse sheaves on such stacks. We will usually reserve the symbols ${\mathscr{A}},{\mathscr{M}},{\mathscr{X}},\ldots$ for stacks. A $1$-morphism between stacks will often be referred to simply as a morphism or a map.
3.1. Euler-Satake characteristic of constructible sheaves on a stack.
Let ${\mathscr{X}}$ be a DM stack of finite type over a field $k$, and ${\rm D}^{b}_{c}({\mathscr{X}},\Lambda)$ denote the derived category of constructible sheaves of $\Lambda$-modules in one of the following settings:
  1. (1)
    $k={\mathbb{C}}$, we take constructible sheaf in the analytic topology, and $\Lambda$ is a field.
  2. (2)
    $k$ is arbitrary, $char(k)=0$ and we consider constructible sheaves in the lisse-etale topos with $\Lambda={\mathbb{Z}}/\ell^{n}{\mathbb{Z}}$ or ${\mathbb{Q}}_{\ell}$.
In any of the aforementioned settings and when $\Lambda$ is a field, one can define the Euler-Satake characteristic $\chi^{orb}({\mathcal{F}})$ of ${\mathcal{F}}\in{\rm D}^{b}_{c}({\mathscr{X}})$. In the analytic setting and where one considers constructible functions, this was defined by Behrend ([5]). And it was considered by several authors for constant coefficients, starting with Satake in the 1950’s. The Euler-Satake characteristic satisfies the following properties and is uniquely determined by these properties:
  1. (i)
    $\chi^{orb}({\mathscr{X}},{\mathcal{F}})\in{\mathbb{Q}}.$
  2. (ii)
    Let ${\mathscr{U}}\subset{\mathscr{X}}$ be open, and ${\mathscr{Z}}$ the closed complement. Then $\chi^{orb}({\mathscr{X}},{\mathcal{F}})=\chi^{orb}({\mathscr{U}},{\mathcal{F}}|_{{\mathscr{U}}})+\chi^{orb}({\mathscr{Z}},{\mathcal{F}}|_{{\mathscr{Z}}})$.
  3. (iii)
    If $f:{\mathscr{X}}\rightarrow{\mathscr{Y}}$ is a finite etale map of degree $d$, then $\chi^{orb}({\mathscr{X}},f^{*}{\mathcal{F}})=d\chi({\mathscr{Y}},{\mathcal{F}})$.
  4. (iv)
    If $X$ is a scheme, the $\chi^{orb}({\mathscr{X}},{\mathcal{F}})=\chi({\mathscr{X}},{\mathcal{F}})$ i.e. the orbifold Euler characteristic is the usual Euler characteristic. Note: In the analytic setting one should instead take the compactly supported Euler characteristic. These are the same in the algebraic setting.
  5. (v)
    Let ${\mathcal{F}}$ be a constructible sheaf on ${\mathscr{X}}$ and ${\mathcal{G}}$ on ${\mathscr{Y}}$. Then $\chi^{orb}({\mathscr{X}}\times{\mathscr{Y}},{\mathcal{F}}\otimes{\mathcal{G}})=\chi^{orb}({\mathscr{X}},{\mathcal{F}})\chi^{orb}({\mathscr{Y}},{\mathcal{G}}).$
  6. (vi)
    Let ${\mathcal{F}}\to{\mathcal{H}}\to{\mathcal{G}}\xrightarrow{[1]}$ be a distinguished triangle in $D_{c}^{b}({\mathscr{X}},\Lambda)$, then $\chi^{orb}({\mathscr{X}},{\mathcal{H}})=\chi^{orb}({\mathscr{X}},{\mathcal{F}})+\chi^{orb}({\mathscr{X}},{\mathcal{G}}).$
  7. (vii)
    One has $\chi({\mathscr{X}},{\mathcal{F}}[d])=(-1)^{d}\chi({\mathscr{X}},{\mathcal{F}})$
Proof.
(Proof of existence) Note that the usual (compactly supported) euler characteristic on schemes or complex analytic spaces satisfies these axioms. This extends uniquely to Deligne-Mumford stacks as follows. If ${\mathscr{X}}$ is a DM stack, then generically it is a quotient by a finite etale group scheme. In particular, one has an open ${\mathscr{U}}\subset{\mathscr{X}}$, a scheme $U$ with an action of a finite etale group scheme $G$ such that ${\mathscr{U}}=[U/G]$. On this open ${\mathscr{U}}$, $\chi({\mathscr{U}},{\mathcal{F}}|_{\mathscr{U}})$ is determined by (iii) above. Now one can proceed via induction on dimension and use property (ii). Alternately, one can also stratify ${\mathscr{X}}$ by quotients of the form above. ∎
We write
$\chi^{orb}({\mathscr{X}})=\chi^{orb}({\mathscr{X}},\Lambda_{X})$
3.2. Perverse sheaves on stacks
In ([20]), Laszlo-Olsson define perverse sheaves in the lisse-étale topology of an Artin stack of finite type (over a fixed field $k$ of $char(k)=0$). We refer to loc. cit. for the details of this construction. Below, we recall some basic facts which will be useful in the following. We denote by $\operatorname{Perv}({\mathscr{X}})$ the corresponding category of perverse sheaves (with coefficients in ${\mathbb{Q}}$).
  1. (1)
    Let $j:{\mathscr{U}}\hookrightarrow{\mathscr{X}}$ be an open immersion. Then $j^{*}$ is $t$-exact.
  2. (2)
    If ${\mathscr{Z}}\hookrightarrow{\mathscr{X}}$ is a closed smooth irreducible substack (or more generally a local complete intersection), then $i_{*}({\mathbb{Q}}[d])$ (with $d=\dim_{{\mathscr{Z}}}$) is a perverse sheaf. Both statements follow directly from the definitions given the analogous statements for schemes, and the fact that local complete intersections are stable under smooth base change.
4. The case of ${}_{N}{\mathscr{A}}_{g}$
Fix $N\geq 3$, and let ${}_{N}{\mathscr{A}}_{g}$ denote the moduli space of principally polarized $g$ dimensional abelian varieties with level $N$ structure over $k$. This is a fine moduli space, and it is nonsingular. Let ${\overline{{}_{N}{\mathscr{A}}_{g}}}$ denote a nonsingular toroidal compactification such that $E={\overline{{}_{N}{\mathscr{A}}_{g}}}-{}_{N}{\mathscr{A}}_{g}$ is a divisor with normal crossings ([10]). In this section, we prove the following theorem:
Theorem 4.1.
Let ${\mathcal{P}}$ be a perverse sheaf on $X={}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ where $N_{i}\geq 3$. Then $\chi(X,{\mathcal{P}})\geq 0$
Proof.
This follows from lemmas 4.3, 4.4 and 4.5 below. ∎
Before proving the theorem, we record the following corollary.
Corollary 4.2.
If $X\subseteq{}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ ($N_{i}\geq 3$) is a smooth subvariety, then
$(-1)^{\dim X}\chi(X)\geq 0$
Lemma 4.3.
Let $A:={}_{N}{\mathscr{A}}_{g}$ and $\bar{A}={\overline{{}_{N}{\mathscr{A}}_{g}}}$. Then $\Omega_{\bar{A}}^{1}(\log E)$ is nef
Proof.
Let $f:U\to A$ denote the universal family of abelian varieties, and let $\overline{F}^{1}$ be the associated Hodge bundle over $\bar{A}$. By [10, chap IV, theorem 7.7],
$\Omega_{\bar{A}}^{1}(\log E)\cong S^{2}\overline{F}^{1}$
Therefore the lemma follows from theorem 2.2 and [21, thm 6.2.12]. ∎
Lemma 4.4.
Let $\bar{A}$ be a smooth projective variety with an snc divisor $E$ such that $\Omega_{\bar{A}}^{1}(\log E)$ is nef. If $P$ is a perverse sheaf on $A=\bar{A}-E$, then $\chi(A,P)\geq 0$.
Proof.
Let $\overline{CC({\mathcal{P}})}\subset T^{*}_{\bar{A}}(\log E)$ be the characteristic cycle defined in section 2.2. This is effective with components given by closures of conormal bundles $T^{*}_{V}A$ for $V\subset A$. The log Dubson-Kashiwara formula of Wu and Zhou (2.1.1) shows that
$\chi(A,P)=\langle\overline{CC(P)},Z\rangle$
where $Z\subset T^{*}_{\bar{A}}(\log E)$ is the zero section. Therefore it is enough to show that this intersection number is nonnegative. Since the log cotangent bundle is nef. Theorem 2.1 yields the desired nonnegativity.
Lemma 4.5.
Let $X_{1}$ and $X_{2}$ be smooth projective varieties with snc divisors $E_{i}$, such that $\Omega_{X_{i}}^{1}(\log E_{i})$ are nef. Then $\Omega_{X_{1}\times X_{2}}^{1}(\log p_{1}^{*}E_{1}+p_{2}^{*}E_{2})$ is nef, where $p_{i}:X_{1}\times X_{2}\to X_{i}$ the projections.
Proof.
This follows from [21, theorem 6.2.12] and the fact that
$\Omega_{X_{1}\times X_{2}}^{1}(\log p_{1}^{*}E_{1}+p_{2}^{*}E_{2})=p_{1}^{*}\Omega_{X_{1}}^{1}(\log E_{1})\oplus p_{2}^{*}\Omega_{X_{2}}^{1}(\log E_{2})$
We conclude the section with some Hodge theoretic variants (which will not be used in the following).
Theorem 4.6.
If $X\subset{}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ is a smooth projective variety, then for all $p$
$(-1)^{\dim X-p}\chi(\Omega_{X}^{p})>0$
Proof.
By [13, 0.3A] $X$ is Kähler hyperbolic, therefore the theorem follows from [13, theorem 2.5]. ∎
Corollary 4.7.
If $X\subset{}_{N_{1}}{\mathscr{A}}_{g_{1}}\times\ldots\times{}_{N_{r}}{\mathscr{A}}_{g_{r}}$ is a smooth projective variety, then
$(-1)^{\dim X}\chi(X)>\dim X$
Proof.
By the Hodge decomposition
$\chi(X)=\sum_{p=0}^{\dim X}(-1)^{p}\chi(\Omega_{X}^{p})$
Remark 4.8.
Since the boundary of the Satake compactification of ${}_{N}{\mathscr{A}}_{g}$ has codimension $g$, it follows that it has plenty of positive dimensional smooth projective subvarieties.
5. Euler characteristics of perverse sheaves on ${\mathscr{A}}_{g,n}$
Let $f:(X,E)\to(Y,D)$ be a projective log smooth map of nc log schemes. Set ${\mathcal{X}}=(X,E)$, ${\mathcal{Y}}=(Y,D)$, $\Omega^{1}_{\mathcal{X}}=\Omega_{X}^{1}(\log E)$ etc. We have an exact sequence
(5.0.1) $0\to f^{*}\Omega^{1}_{\mathcal{Y}}\to\Omega^{1}_{\mathcal{X}}\to\Omega^{1}_{{\mathcal{X}}/{\mathcal{Y}}}\to 0$
Lemma 5.1.
In the above situation, suppose that $\Omega^{1}_{\mathcal{Y}}$ is nef and that $\Omega_{{\mathcal{X}}/{\mathcal{Y}}}^{1}=f^{*}V$, where $V$ is nef. Then $\Omega_{\mathcal{X}}^{1}$ is nef.
Proof.
This follows from the fact that the nef property is closed under pull backs and extensions [21, thm 6.2.12].
Fix $N\geq 3$. Since ${}_{N}{\mathscr{A}}_{g}$ is a fine moduli space, we have a universal abelian scheme ${}_{N}{\mathscr{A}}_{g,1}\to{}_{N}{\mathscr{A}}_{g}$. Let ${}_{N}{\mathscr{A}}_{n}={}_{N}{\mathscr{A}}_{1}\times_{{}_{N}{\mathscr{A}}_{g}}\ldots\times_{{}_{N}{\mathscr{A}}_{g}}{}_{N}{\mathscr{A}}_{1}$ denote the $n$-fold fibre product. This a fine moduli space, whose $S$-valued points are $g$ dimensional abelian schemes of $S$ with a principal polarization, level $N$ structure, and $n$ sections.
Theorem 5.2.
If ${\mathcal{P}}$ be a perverse sheaf on $X={}_{N}{\mathscr{A}}_{g_{1},n_{1}}\times\ldots\times{}_{N}{\mathscr{A}}_{g_{r},n_{r}}$, then $\chi(X,{\mathcal{P}})\geq 0$.
Proof.
Let
$\overline{X}=\overline{{}_{N}{\mathscr{A}}_{g_{1},n_{1}}}\times\ldots\times\overline{{}_{N}{\mathscr{A}}_{g_{r},n_{r}}}$
be a product of toroidal compactifications. By a theorem of Faltings-Chai [10, chap VI, thm 1.1], we can find an nc log scheme $(P,D)$ with morphism $f:(P,D)\to\overline{X}$ such that
  1. (1)
    $P-D=X$ and $f|_{P-D}$ coincides with the product of canonical maps ${}_{N}{\mathscr{A}}_{g_{i},n}\to{}_{N}{\mathscr{A}}_{g_{i}}$.
  2. (2)
    $f$ is toroidal [10, chap VI, thm 1.13], and therefore log smooth by [16, prop 3.4].
  3. (3)
    $\Omega^{1}_{(P,D)/(\overline{X},E)}$ is the pull back of a Hodge bundle on $\overline{X}$.
The theorem now follows from lemmas 4.4 and 5.1, and theorem 2.2.
Theorem 5.3.
If $P$ is a perverse sheaf on ${\mathscr{X}}={\mathscr{A}}_{g_{1},n_{1}}\times\ldots\times{\mathscr{A}}_{g_{r},n_{r}}$, then
$\chi^{orb}({\mathscr{X}},P)\geq 0$
In particular, if ${\mathscr{Y}}\subset{\mathscr{X}}$ is a smooth closed irreducible substack, then
$(-1)^{\dim{\mathscr{Y}}}\chi^{orb}({\mathscr{Y}})\geq 0$
Furthermore, we have strict inequality
$(-1)^{\dim{\mathscr{Y}}}\chi^{orb}({\mathscr{Y}})>0$
if ${\mathscr{Y}}\subset{\mathscr{A}}_{g}$ is proper.
Proof.
Since ${\mathscr{X}}=\prod[{}_{N}{\mathscr{A}}_{g_{i},n_{i}}/PSp_{2g_{i}}({\mathbb{Z}}/N{\mathbb{Z}})]$, this follows from 3.1 (iii) and theorems 4.1 and 4.6.
Remark 5.4.
The theorem cannot be improved to a strict inequality in general, because if $n>0$, then $\chi^{orb}({\mathscr{A}}_{g,n})=0$. However, we do not know if there is an example of $X\subset{\mathscr{A}}_{g}$ with $\chi(X)=0$.
Recall that a morphism $f:X\to Y$ is called semismall if for all $k$
$2k\leq\dim Y-\dim\{y\in Y\mid\dim f^{-1}(y)=k\}$
Corollary 5.5.
If $\pi:X\to{\mathscr{A}}_{g,n}$ is a semismall proper map from a smooth irreducible stack, then
$(-1)^{\dim X}\chi^{orb}(X)\geq 0$
Proof.
By [9, prop 4.2.1] $\pi_{*}{\mathbb{Q}}[\dim X]$ is perverse. ∎
We will give a concrete application of this below in Proposition 8.2.
6. Key Inequality
In this section, we shall work in the setting of varieties over an algebraically closed field $k$ of characteristic 00. Given a variety $X$, we say that $X$ has the perverse Euler characteristic, or $PEC$-property, or simply that $X$ is $PEC$, if for every perverse sheaf ${\mathcal{K}}\in\operatorname{Perv}(X)$, $\chi(X,{\mathcal{K}})\geq 0$. If $X$ comes equipped with a ${\mathbb{G}}_{m}$-action, we will use similar terminology in the setting of monodromic perverse sheaves: $X$ has the $PEC_{m}$ property if for all ${\mathcal{K}}\in Perv_{m}(X)$, $\chi(X,{\mathcal{K}})\geq 0$. We shall say that $X$ has the $PEC[d]$ (resp. $PEC_{m}[d]$)-property if $\chi(X,{\mathcal{K}}[d])\geq 0$ for all ${\mathcal{K}}\in\operatorname{Perv}(X)$ (resp. $\operatorname{Perv}_{m}(X))$. Clearly $PEC[d]$ is the same as $PEC$ if $d$ is even, and $PEC[-1]$ otherwise.
[…]
Theorem 6.1.
Let $(X,Y,f,U,Z)$ be an admissible datum as above. Suppose that $Y$ and $Z$ have the $PEC[d]$-property. Then $X$ also has the $PEC[d]$-property. The analogous assertion holds in the monodromic setting, i.e. when the datum admits a ${\mathbb{G}}_{m}$-action and $Y$ and $Z$ have the $PEC_{m}[d]$-property, then $X$ has the $PEC_{m}[d]$-property.
We shall prove the theorem via induction on dimension, and henceforth assume that it holds for admissible data $(X^{\prime},Y^{\prime},f^{\prime},U^{\prime},Z^{\prime})$ with $\dim(Y^{\prime})<\dim Y=n$.
6.1. Preliminary reduction
In this subsection, we assume the existence of a non-constant regular function $h:Y\to{\mathbb{A}}^{1}$ such that $Z=g^{-1}(0)$, where $g=h\circ f$. Let $D=f(Z)=h^{-1}(0)$. We prove Theorem 6.1 under these conditions.
Lemma 6.2.
Let $K\in\operatorname{Perv}(X)$, and let $(K_{U},K_{Z},c,v)$ be the corresponding object in $B_{g}(U,Z)$ (see theorem 2.6). Then
$\chi(K)=\chi(\tilde{j}_{*}K_{U})+\chi(K_{Z})-\chi(\psi_{g}^{u}K_{U})$
Similarly, for $K\in\operatorname{Perv}(Y)$, with the obvious notation we have
$\chi(K)=\chi(j_{*}K_{U})+\chi(K_{D})-\chi(\psi_{h}^{u}K_{U})$
Proof.
From (2.3.2) and (2.3.1), we obtain
$\begin{split}\chi(K)&=\chi(\ker(v,-\beta))-\chi(\psi_{g}^{u}K_{U})\\ &=\chi(\Xi_{g}K_{U})+\chi(K_{Z})-2\chi(\psi_{g}^{u}K_{U})\\ &=\chi(\tilde{j}_{*}K_{U})+\chi(K_{Z})-\chi(\psi_{g}^{u}K_{U})\end{split}$
The proof of the second part is similar.
Lemma 6.3.
  1. (1)
    Given an admissible datum $(X,Y,f,U,Z)$ such that $Y$ has the $PEC$-property. If $K_{U}\in\operatorname{Perv}(U)$, then
    $\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U})\geq 0.$
  2. (2)
    Suppose the $(X,Y,f,U,Z)$ come equipped with compatible ${\mathbb{G}}_{m}$ actions, and that $Y$ satisfies the $PEC_{m}[d]$-property. If $K_{U}\in\operatorname{Perv}_{m}(U)$ then
    $(-1)^{d}(\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U}))\geq 0.$
Proof.
We prove the first part, the proof of the second part is the same. Without loss of generality, we can assume that the ground field $k={\mathbb{C}}$. Given $a\in{\mathbb{N}}$, let $L_{a}$ be pull back to $U$ of the local system on ${\mathbb{C}}^{*}$ with monodromy given by the $a\times a$ Jordan block
$\begin{pmatrix}1&1&0&0&\ldots\\ 0&1&1&0&\ldots\\ 0&0&1&1&\ldots\\ &&\ldots&&\end{pmatrix}$
We choose (and fix) $a\gg 0$. We may replace $K_{U}$ by $K_{U}\otimes L_{a}$, since the expression $\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U})$ just gets multiplied by $\operatorname{rank}L_{a}$. It follows from [24, prop 3.1], that the monodromy of $\psi_{g}^{u}K_{U}$ and $\psi_{h}^{u}K_{U}$ is now trivial. Let $M=j_{!*}K_{U}$. and $\tilde{M}=\tilde{j}_{!*}K_{U}$. By [27, prop 4.7], $\tilde{M}$ corresponds to
$(\tilde{M}_{U}=K_{U},0,0,0)\in B_{g}(U,Z)$
and $M$ corresponds to
$(M_{U}=K_{U},0,0,0)\in B_{h}(U,D$
We have
$j_{*}K_{U}\cong f_{*}\tilde{j}_{*}K_{U}$
by proper base change. Also
$\psi_{h}K_{U}\cong f_{*}\psi_{g}K_{U}$
by [1, exp XIII (2.1.7.1)]. This isomorphism is compatible with monodromy. It follows that this isomorphism restricts to the unipotent parts. Therefore
$\begin{split}\chi(\tilde{j}_{*}K_{U})-\chi(\psi_{g}^{u}K_{U})&=\chi(j_{*}K_{U})-\chi(\psi_{h}^{u}K_{U})\\ &=\chi(M)\\ &\geq 0\end{split}$
The last inequality follows from the assumptions since $M\in\operatorname{Perv}(Y)$.
Corollary 6.4.
Consider an admissible datum $(X,Y,f,U,Z)$ as above. If $Y$ and $Z$ satisfy the $PEC$-property, then so does $X$. In the presence of ${\mathbb{G}}_{m}$-actions, analogous statement holds if $Y$ and $Z$ have the $PEC_{m}[d]$-property.
6.2. Second reduction.
In this subsection, we prove Theorem 6.1 in the setting where we have a function $h:Y\rightarrow{\mathbb{A}}^{1}$ such that $f(Z)\subseteq h^{-1}(0).$ We shall make this assumption for the remainder of this subsection.
Lemma 6.5.
Let $Z$ be a Cartier divisor on a variety $X$. Suppose $Z=Z_{1}\cup Z_{2}$ where each $Z_{i}$ is a union of components of $Z$ (with non components in common). If the $PEC$-property holds for each $Z_{i}$, then it also holds for all objects of $Perv(Z)$. The analogous assertion holds in the monodromic setting.
Proof.
Let $P\in\operatorname{Perv}(Z,S)$. Then $\chi(Z,P)=\chi(Z_{1},P)+\chi(Z_{2},j_{!}P)$, where $j:Z_{2}-Z_{1}\to Z_{2}$ is inclusion. On the other hand, under the given hypotheses $Z_{1}\cap Z_{2}$ is an effective Cartier divisor on $Z_{2}$, and it equals $Z_{2}\setminus(Z_{2}\setminus Z_{1})$. It follows that $j_{!}P$ is perverse, and the result now follows. In the monodromic setting, we note that the same argument works because $j_{!}$ preserves monodromic perverse sheaves. ∎
Lemma 6.6.
  1. (1)
    With notation and hypotheses as in theorem 6.1, suppose furthermore that there exists a function $h:Y\rightarrow{\mathbb{A}}^{1}$ such that $h^{-1}(0)\supseteq f(Z)$. Then the conclusion of the theorem holds.
  2. (2)
    Suppose that the tuple $(X,Y,f,U,Z)$ and $h$ as above all come equipped with a ${\mathbb{G}}_{m}$-equivariant action. If $Y$, $Z$ satisfy the $PEC_{m}[d]$ property, then so does $X$.
Proof.
For (1), let $D_{Y}:=h^{-1}(0)$ and $D_{X}:=g^{-1}(0)$ where $g:=h\circ f:X\rightarrow{\mathbb{A}}^{1}$. Note that $f(D_{X})=f(f^{-1}(D_{Y}))=D_{Y}$. Let $V=Y\setminus D_{Y}\subset U$. It follows that this gives rise to an admissible datum $(X,Y,f,D_{X},V)$. Moreover, we claim that for every perverse sheaf $K\in\operatorname{Perv}(D_{X})$, $\chi(D_{X},K)\geq 0$. In order to see this, first note that the analogous assertion holds for $D_{Y}$ since it holds for $Y$ (and push-forward from a closed subscheme is exact for the perverse t-structure). We may write $D_{X}=D^{\prime}_{X}\cup Z$ as a union of Cartier divisors. Similarly, set $D_{Y}=D^{\prime}_{Y}\cup f(Z)$. Restricting $f$ to $D^{\prime}_{X}$ gives rise to an admissible tuple $(D^{\prime}_{X},D^{\prime}_{Y},f,Z\cap D^{\prime}_{X},U\setminus V)$. We may assume by induction on dimension of $Y$ (or equivalently $X$) that the Euler characteristic of every perverse sheaf on $D^{\prime}_{X}$ is non-negative. Note that this property holds for both $D^{\prime}_{Y}$ and $Z\cap D^{\prime}_{X}$ since it holds for $Y$ and $Z$. Now by Lemma 6.5 above, we may conclude that the Euler characteristic of every perverse sheaf on $D_{X}$ is non-negative, as claimed. Now for the tuple $(X,Y,f,V,D_{X})$, the result follows from Corollary 6.4. Therefore, this proves the desired property for all perverse sheaves on $X$.
In the monodromic setting of (2), the same proof goes through.
6.3. The general case
In order to complete the proof, we use deformation to the normal cone. Given a pair $Z\subset X$ with $Z$ a closed subvariety, let $C(Z/X)$ denote the normal cone of $Z$ in $X$, and let $C^{*}(Z/X):=C(Z/X)\setminus Z$, $U:=X\setminus Z$. The deformation to the normal cone construction gives rise to a diagram:
[…]
We remind the reader that $V(Z/X)$ is obtained as follows. First, consider the blow-up $B(Z/X):=Bl_{Z\times\{0\}}(X\times{\mathbb{A}}^{1})$. Then set $V(Z/X):=B(X/Z)\setminus(Bl_{Z}(X))$ and $V^{*}(Z/X)=V(Z/X)-Z\times{\mathbb{A}}^{1}.$ Algebraically, one obtains $V(Z/X)$ as follows. Let ${\mathcal{I}}\subset{\mathcal{O}}_{X}$ denote the ideal of $Z$. Then $V(Z/X)={\bf Spec}(R_{Z})$, where $R_{Z}:=\bigoplus_{n\in{\mathbb{Z}}}{\mathcal{I}}^{n}t^{-n}\subset{\mathcal{O}}_{X}[t,t^{-1}]$ is the (extended) Rees algebra. We see from this that the natural projection map $p:V(Z/X)\rightarrow X$ is flat. If $Z$ is a Cartier divisor, then all fibers are smooth of dimension 1, and hence $p$ is smooth. We note that the bottom three rows of the previous diagram come equipped with ${\mathbb{G}}_{m}$-actions, and the morphisms in the diagram respect these actions. These constructions are functorial in the following sense.
Lemma 6.7.
If $Z\subset X$ and $Z^{\prime}\subset X^{\prime}$ are closed subvarieties, and if $f:X\to X^{\prime}$ is a morphism such that $Z=f^{-1}Z^{\prime}$, then there is a ${\mathbb{G}}_{m}$-equivariant morphism $F:V(Z/X)\to V(Z^{\prime}/X^{\prime})$ which induces a map of the corresponding diagrams above. If $f$ is proper, then $F$ is proper.
Proof.
Suppose $Z=f^{-1}Z^{\prime}$, then $f$ induces a map $F_{1}:B(Z/X)\to B(Z^{\prime}/X^{\prime})$ between the blow ups such that $F_{1}^{-1}(Bl_{Z^{\prime}}X^{\prime}\times\{0\})=Bl_{Z}X\times\{0\}$. This restricts to give $F:V(Z/X)\to V(Z^{\prime}/X^{\prime})$. If $f$ is proper, then $F_{1}$ is proper, so the same is true for $F$. ∎
We also need the following.
Lemma 6.8.
Let $Z\subset X$ be an effective Cartier divisor. Then $X$ has the $PEC$-property if and only if $V(Z/X)$ has the $PEC_{m}[-1]$-property.
Proof.
Since $Z\hookrightarrow X$ is a Cartier divisor, then $p:V(Z/X)\rightarrow X$ is smooth of relative dimension 1. It follows that if ${\mathcal{K}}\in\operatorname{Perv}(X)$, then $p^{*}({\mathcal{K}})[1]$ is a monodromic perverse sheaf on $V(Z/X)$. Moreover, the adjunction ${\mathcal{K}}\rightarrow Rp_{*}p^{*}{\mathcal{K}}$ is an isomorphism. The lemma easily follows. ∎
Proof of Theorem 6.1.
Suppose we are given an admissible datum $(X,Y,f,U,Z)$ such that $Y$ and $Z$ have the $PEC$ property. Since $Y$ is quasiprojective, we can choose a Cartier divisor $D\supseteq f(Z)$. Let us write $f^{-1}D=Z\cup Z^{\prime}$, where $Z^{\prime}$ is the union of components $Z_{i}^{\prime}$ not contained in $Z$. Then $(Z_{i}^{\prime},f(Z_{i}^{\prime}),f|_{Z_{i}^{\prime}},U\cap Z_{i}^{\prime},Z_{i}^{\prime}\cap Z)$ are admissible data satisfying the assumptions of the theorem. By induction and Lemma 6.5, we find that $Z^{\prime}$ has the $PEC$ property. Since $Z$ also has the $PEC$ property, we can conclude that so does $Z\cup Z^{\prime}$ by Lemma 6.5. Therefore $C(Z\cup Z^{\prime}/X)$ has the $PEC_{m}[-1]$ property by Lemma 6.8. This Lemma also shows that $V_{Y}:=V(D/Y)$ has the $PEC_{m}[-1]$-property. Let $V_{X}=V(Z\cup Z^{\prime}/X)$. The map $f$ induces a proper map $F:V_{X}\to V_{Y}$ by Lemma 6.7. Therefore we have admissible datum $(V_{X},V_{Y},F,U\times{\mathbb{G}}_{m},Z^{\prime\prime})$, where $Z^{\prime\prime}=C(Z\cup Z^{\prime}/X)\cup(Z\cup Z^{\prime})\times{\mathbb{A}}^{1}$. Lemma 6.5 shows that $Z^{\prime\prime}$ has the $PEC_{m}[-1]$ property. Moreover, we have $h^{-1}(0)\supseteq F(Z^{\prime\prime})$, where $h:V_{Y}\rightarrow{\mathbb{A}}^{1}$ is the projection. In particular, we are in the setting of Lemma 6.6. It follows that $V_{X}$ has the $PEC_{m}[-1]$-property, and therefore $X$ has the $PEC$-property by Lemma 6.8.
7. Extended Torelli map
Let $\overline{{\mathscr{M}}_{g}}$ (respectively ${\mathscr{M}}_{g}$) denote the stack parameterizing stable (respectively smooth) curves of genus $g$ [9]. Recall that a stable curve $C$ is of compact type if its dual graph is a tree, or equivalently if the identity component of its Picard scheme $\operatorname{Pic}^{0}(C)$ is an abelian variety. Let
${\mathscr{M}}_{g}\subset{\mathscr{M}}_{g}^{c}\subset{\overline{{\mathscr{M}}_{g}}}$
be the open subset parameterizing curves of compact type.
By a theorem of Mumford and Namikawa [26, §9D] the Torelli map $\tau:{\mathscr{M}}_{g}\to{\mathscr{A}}_{g}$ extends to a map $\tau:\overline{{\mathscr{M}}_{g}}\to\overline{{\mathscr{A}}_{g}}$ to the toroidal compactification associated to the second Voronoi decomposition. Alexeev [2, 3] gave a modular interpretation of both the target and the map. On ${\mathscr{M}}_{g}^{c}$, $\tau$ simplifies to
[…]
where $C_{i}$ are the irreducible components of $C$. An immediate consequence of properness of $\overline{{\mathscr{M}}_{g}}$ and the theorem of Mumford-Namikawa is:
Lemma 7.1.
The map $\tau$ restricts to a proper map $\tau:{\mathscr{M}}_{g}^{c}\to{\mathscr{A}}_{g}$. Therefore $\tau({\mathscr{M}}^{c}_{g})$ is the closure of $\tau({\mathscr{M}}_{g})$ in ${\mathscr{A}}_{g}$
We will give a more direct proof of properness below. Given $N>0$, let ${}_{N}{\mathscr{M}}_{g}$ denote the stack of smooth curves with level $N$ structure. Let $\overline{{}_{N}{\mathscr{M}}_{g}}$ (respectively ${}_{N}{\mathscr{M}}^{c}_{g}$) denote the normalization of $\overline{{\mathscr{M}}_{g}}$ (respectively ${\mathscr{M}}^{c}_{g}$) in the function field of ${}_{N}{\mathscr{M}}_{g}$. The stack ${{}_{N}{\mathscr{M}}^{c}_{g}}$ can be identified with the moduli stack of stable curves of compact type with level $N$ structure (by contrast $\overline{{}_{N}{\mathscr{M}}_{g}}$ does not have a good modular interpretation). Thus we can lift $\tau$ to a map $\tau:{}_{N}{\mathscr{M}}^{c}_{g}\to{}_{N}{\mathscr{A}}_{g}$.
Let $\overline{{\mathscr{M}}_{g,n}}$ be the Deligne-Mumford stack of genus $g$ stable curves with $n$ base points [19]. A geometric point of this space stack corresponds a semistable curve $(C,p_{1},\ldots,p_{n})$ of arithmetic genus $g$ with $n$ smooth base points such that the total number of base and double points on any smooth rational component is at least three. We refer to this is as an $n$-pointed stable curve. When $n>0$, there exists a morphism $\pi_{n}:\overline{{\mathscr{M}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n-1}}$ called a contraction. For a curve without rational components, $\pi_{n}$ simply forgets $p_{n}$. Given integers $g=g_{1}+g_{2}$, and a partition of $\{1,\ldots,n\}$ into two subsets of cardinality $n_{1}$ and $n_{2}$, Knudsen also constructs a gluing map
$gl:\overline{{\mathscr{M}}_{g_{1},n_{1}+1}}\times\overline{{\mathscr{M}}_{g_{2},n_{2}+1}}\to\overline{{\mathscr{M}}_{g,n}}$
which takes $((C,p_{1},\ldots,p_{n_{1}+1}),(C^{\prime},p^{\prime}_{1},\ldots,p^{\prime}_{n_{2}+1}))$ to the curve $C\coprod C^{\prime}/(p_{n_{1}+1}\sim p^{\prime}_{n_{2}+1})$ with the remaining base points $p_{1},\ldots,p_{n_{1}},p_{1}^{\prime},\ldots,p_{n_{2}}^{\prime}$ enumerated appropriately.
We recall two basic facts regarding these constructions (see [19]).
  1. (1)
    First, the contraction map is equivalent to the universal curve over ${\mathscr{M}}_{g,n-1}$ (see Lemma 7.2).
  2. (2)
    Second, the map $gl$ is finite.
The forgetful map $\overline{{\mathscr{M}}_{g,n}}\to\overline{{\mathscr{M}}_{g}}$ is the composition of contractions. Let ${{\mathscr{M}}_{g,n}}$ and ${{\mathscr{M}}^{c}_{g,n}}$ denote the preimages of ${\mathscr{M}}_{g}$ and ${\mathscr{M}}_{g}^{c}$ respectively. We note that ${\mathscr{M}}^{c}_{1,n}={\mathscr{M}}_{1,n}$. When $g>1$, the complement ${\mathscr{M}}^{c}_{g,n}-{\mathscr{M}}_{g,n}$ is a union of divisors, each of which can be described as the image of a gluing map with respect to a partition $g=g_{1}+g_{2}$ with $g_{i}>0$.
Lemma 7.2.
The contraction map ${\mathscr{M}}^{c}_{g,n+1}\to{\mathscr{M}}^{c}_{g,n}$ is equivalent to the universal curve over ${\mathscr{M}}^{c}_{g,n}$.
Proof.
Given a scheme $S$, an $S$-valued point of $\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{M}}^{c}_{g,n+1}$) is a stable curve ${\mathcal{C}}\to S$ with $n$ sections $\sigma_{1},\ldots,\sigma_{n}$, such that the geometric fibres are $n$-pointed genus $g$ stable curves (of compact type). The universal curve $\overline{{\mathscr{Z}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{Z}}^{c}_{g,n}\to{\mathscr{M}}^{c}_{g,n}$) is the stack whose $S$-points are stable curves ${\mathcal{C}}\to S$ with $n+1$ sections $\sigma_{1},\ldots,\sigma_{n+1}$, such that $({\mathcal{C}}\to S,\sigma_{1},\ldots,\sigma_{n})$ defines a point of $\overline{{\mathscr{M}}_{g,n}}$ (respectively ${\mathscr{M}}^{c}_{g,n}$). There is an obvious functor $\overline{{\mathscr{M}}_{g,n+1}}\to\overline{{\mathscr{Z}}_{g,n}}$, which sends to ${\mathscr{M}}^{c}_{g,n+1}$ to ${\mathscr{Z}}^{c}_{g,n}$. Knudsen [19, §2] defines an inverse $s:\overline{{\mathscr{Z}}_{g,n}}\to\overline{{\mathscr{M}}_{g,n+1}}$ called stabilization. When $S=\operatorname{Spec}k$, then roughly
$s({\mathscr{C}},\sigma_{1},\ldots,\sigma_{n+1})=\begin{cases}({\mathcal{C}},\sigma_{1},\ldots,\sigma_{n+1})&\text{if this is an $(n+1)$-pointed stable curve}\\ ({\mathcal{C}}^{\prime},\sigma_{1}^{\prime},\ldots,\sigma_{n+1}^{\prime})&\text{if $\sigma_{n+1}$ is a singular point or another section}\\ \end{cases}$
where ${\mathcal{C}}^{\prime}$ is obtained by adding a ${\mathbb{P}}^{1}$, such that on the dual graph it is inserted either as a pendant vertex or on an edge. Since this operation takes trees to trees, it follows that $s({\mathscr{Z}}^{c}_{g,n})\subset{{\mathscr{M}}^{c}_{g,n+1}}$. Thus ${\mathscr{M}}^{c}_{g,n+1}$ is equivalent to ${\mathscr{Z}}^{c}_{g,n}$. ∎
Proposition 7.3.
The morphism $\tau:{\mathscr{M}}_{g,n}\to{{\mathscr{A}}_{g,n}}$ given on geometric points by
$\tau[(C,p_{1},\ldots,p_{n})]=[(\operatorname{Pic}^{0}(C),(2g-2)p_{1}-K,\ldots,(2g-2)p_{n}-K)],$
where $K$ is the canonical divisor class, is well defined and injective on points. It extends to a proper morphism (denoted by same letter) $\tau:{\mathscr{M}}^{c}_{g,n}\to{{\mathscr{A}}_{g,n}}$. Furthermore, this lifts to a morphism $\tau:{}_{N}{\mathscr{M}}^{c}_{g,n}\to{}_{N}{{\mathscr{A}}_{g,n}}$ for any $N$, where ${}_{N}{\mathscr{M}}^{c}_{g,n}={}_{N}{\mathscr{M}}^{c}_{g}\times_{{\mathscr{M}}^{c}_{g}}{\mathscr{M}}^{c}_{g,n}$
Proof.
For existence, it suffices to treat the case where $N\geq 3$, since the other cases can be handled by taking quotients. Then $M={}_{N}{\mathscr{M}}^{c}_{g,n}$ is fine with a universal family ${\mathcal{C}}={}_{N}{\mathscr{Z}}_{g,n}$ over it. The curve ${\mathcal{C}}$ carries universal sections $\sigma_{1},\ldots,\sigma_{n}$. The relative Picard algebraic space $\operatorname{Pic}({\mathcal{C}}/M)$ exists by [8, chap 9.4, thm 1]. Furthermore, this is a disjoint union of subschemes $P^{d}=\operatorname{Pic}^{d}({\mathcal{C}}/M)$ parameterizing line bundles with degree (defined as $\chi(L)-g+1$) $d$ on the fibres. Note that $P^{0}\to M$ is a $g$-dimensional abelian scheme with level $N$-structure, and $P^{d}$ is a torsor over $P^{0}$. When $d=2g-2$, there is a canonical section given by the class of $\omega_{{\mathcal{C}}/M}$, which allows us to identify $P^{2g-2}$ with $P^{0}$. The class of $(2g-2)\sigma_{i}$ defines a section of $\operatorname{Pic}({\mathcal{C}}/M)$. In fact, by restricting to ${}_{N}{\mathscr{M}}_{g,n}$, we can see that $(2g-2)\sigma_{i}$ lies in $P^{2g-2}$. We can view these as sections of $P^{0}$ by the above identification. Therefore, the universal property of ${}_{N}{\mathscr{A}}_{g,n}$ gives map $\tau:M\to{}_{N}{{\mathscr{A}}_{g,n}}$. On points of $M$, it is clearly described (with some abuse of notation) as
$\tau[(C,\sigma_{1},\ldots,\sigma_{n})]=[(\operatorname{Pic}^{0}(C),(2g-2)\sigma_{1}-\omega_{C},\ldots,(2g-2)\sigma_{n}-\omega_{C})],$
Abel’s theorem plus Torelli’s theorem shows that $\tau$ is injective on ${\mathscr{M}}_{g,n}$.
To check properness, we use the valuative criterion. Let $R$ be a discrete valuation ring with generic point $\eta$, and suppose that we are given a $2$-commutative diagram
[…]
with solid arrows. Then after replacing $R$ by a finite extension, we will construct an arrow $h$ making the new diagram $2$-commute. Since $\overline{{\mathscr{M}}_{g,n}}$ is proper [19], we do get an arrow $f:\operatorname{Spec}R\to\overline{{\mathscr{M}}_{g,n}}$ (after a finite extension), making the resulting diagram commute. In geometric language, we are given a $g$-dimensional principally polarized abelian scheme ${\mathcal{A}}\to\operatorname{Spec}R$, with $n$ sections $\alpha_{i}$. The map $f$ gives a stable $n$-pointed genus $g$ curve $({\mathcal{C}}\to\operatorname{Spec}R,\sigma_{i})$ such that
$({\mathcal{A}},\alpha_{i})_{\eta}\cong\tau({\mathcal{C}}_{\eta},\sigma_{i,\eta})=(\operatorname{Pic}^{0}({\mathcal{C}}/\operatorname{Spec}R),(2g-2)\sigma_{i}-\omega_{\mathcal{C}})_{\eta}$
By the theory of Néron models [8, chap 9.5, thm 4] , we have ${\mathcal{A}}\cong\operatorname{Pic}^{0}({\mathcal{C}}/\operatorname{Spec}R)$. It follows that ${\mathcal{C}}$ is of compact type. Since ${\mathcal{A}}/\operatorname{Spec}R$ is separated, we have an isomorphism
$({\mathcal{A}},\alpha_{i})\cong(\operatorname{Pic}^{0}({\mathcal{C}}/\operatorname{Spec}R),(2g-2)\sigma_{i}-\omega_{\mathcal{C}})$
Therefore the image of $f$ lies in ${\mathscr{M}}^{c}_{g,n}$, and moreover $h=f$ makes the diagram commute. The extension is to the level $N$ case is straightforward and left to the reader. ∎
8. Euler characteristics of perverse sheaves on ${\mathscr{M}}_{g,n}$
Let $j:{\mathscr{M}}_{g}\subset\tau({\mathscr{M}}_{g}^{c})=\overline{\tau({\mathscr{M}}_{g})}$ denote the inclusion, where $\tau$ is discussed in the previous section. We start with a warm up to the main result.
Proposition 8.1.
If $P$ is a perverse sheaf on ${\mathscr{M}}_{g}$, then $\chi(j_{!*}P)\geq 0$.
Proof.
This follows from theorem 5.3. ∎
When $g=2$, the map $\tau$ is finite and surjective, and therefore semismall. So corollary 5.5 applies. When $g>2$, the map is not finite, however:
Proposition 8.2.
If $g=3$ then $\tau$ is semismall. Therefore $\chi({\mathscr{M}}_{3}^{c})\geq 0$
Proof.
When $g=3$, we claim that the fibres of $\tau$ are at most one dimensional, and these lie over the codimension $2$ subset of ${\mathscr{A}}_{3}$ parameterizing reducible abelian varieties. To prove the claim, observe that a singular curve $C$ in ${\mathscr{M}}_{3}^{c}$ is either
  1. (a)
    a union $C=C_{1}\cup C_{2}$ of an genus $1$ curve and a genus $2$ curve meeting at a point $p$,
  2. (b)
    a union $C=C_{1}\cup C_{2}\cup C_{2}$ of three genus $1$ curves where the dual graph is chain,
  3. (c)
    or a union $C=C_{0}\cup C_{1}\cup C_{2}\cup C_{3}$, where $C_{1},C_{2},C_{3}$ have genus $1$, $C_{0}={\mathbb{P}}^{1}$, and the dual graph is a trivalent tree with $C_{0}$ as the root.
In case (a), $C_{1}$ can be taken to be an elliptic curve with $p$ as its origin. Since $\tau([C])=[\operatorname{Pic}^{0}(C_{1})\times\operatorname{Pic}^{0}(C_{2})]$, the map forgets $p\in C_{2}$. Therefore the fibre $\tau^{-1}(\tau([C]))$ can be identified with $C_{2}$. In case (b), we take $C_{1}\cap C_{2}$ as the origins of $C_{1}$, $C_{2}$, and $C_{2}\cap C_{3}$ as the origin of $C_{3}$. The point $p\in C_{2}\cap C_{3}$ is distinct from 00 in this case; we treat (c) as the limiting case where $p\to 0$. Then $\tau^{-1}(\tau([C]))$ with $C_{2}$. The claim follows from this, and it implies that $\tau$ is also semismall, when $g=3$. The inequality follows from corollary 5.5. ∎
Remark 8.3.
We will prove the corresponding inequality for all $g$. However, the map ${\mathscr{M}}_{g}^{c}\to{\mathscr{A}}_{g}$ is not semismall when $g>3$, so it requires a different strategy.
Theorem 8.4.
Let ${\mathscr{M}}={\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{c}$, where either $g_{i}>1$ or $g_{i}=1$ and $n_{i}>0$. If ${\mathcal{K}}$ is perverse sheaf on ${\mathscr{M}}$, then
$\chi^{orb}({\mathscr{M}},{\mathcal{K}})\geq 0$
Proof.
There is no loss in assuming that $g_{1}\geq g_{2}\geq\ldots\geq 0$. Let $<_{L}$ denote the lexicographic order on the set of the vectors $(g_{1},g_{2},\ldots)$. For example,
$(1,1,1)<_{L}(2,1)<_{L}(3)$
This is a well ordering, so we can use it to do induction. Fix $N\geq 3$, and let $X={}_{N}{\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{}_{N}{\mathscr{M}}_{g_{r},n_{r}}^{c}$. It is enough to prove that
(8.4.1) $\chi(X,K)\geq 0$
for $K\in\operatorname{Perv}(X)$. Suppose that $g_{1}=1$, i.e. all $g_{i}=1$. Then we have an affine open immersion $j:X\to\prod{}_{N}{\mathscr{A}}_{1,n_{i}}$. Therefore $j_{*}K$ is perverse. Consequently theorem 5.2 implies
$\chi(X,K)=\chi(\prod{}_{N}{\mathscr{A}}_{1,n_{i}},j_{*}K)\geq 0$
Therefore we may assume that $g_{1}>1$. We prove (8.4.1) by induction on $(g_{1},g_{2},\ldots)$. For each $g_{i}\geq 2$ and decompositions $g_{i}=g_{1}^{\prime}+g_{2}^{\prime}$, with $g_{i}^{\prime}>0$ and $n_{i}=n_{1}^{\prime}+n_{2}^{\prime}$,, we have a map
$id\times\ldots gl\ldots:{}_{N}{\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times({}_{N}{\mathscr{M}}_{g_{1}^{\prime},n_{1}^{\prime}}^{c}\times{}_{N}{\mathscr{M}}_{g_{2}^{\prime},n_{2}^{\prime}}^{c})\times\ldots\to X$
Let $Z=X-\prod{}_{N}{\mathscr{M}}_{g_{i},n_{i}}$. When $g_{1}>1$, $Z$ is a union of images of the above gluing maps.
We claim that $\chi(Z,K^{\prime\prime})\geq 0$ for $K^{\prime\prime}\in\operatorname{Perv}(Z)$. Since $\operatorname{Perv}(Z)$ is Artinian, we can assume that $K^{\prime\prime}$ is simple. Then $K^{\prime\prime}$ is the image of a perverse sheaf $L$ in some $\prod{}_{N}{\mathscr{M}}_{g_{i}^{\prime},n_{i}^{\prime}}$, with $(g_{1}^{\prime},g_{2}^{\prime},\ldots)<_{L}(g_{1},g_{2},\ldots)$. Therefore the claim follows by induction.
By taking products of $\tau$, we get a morphism that we also denote by the same letter
$\tau:X\to\prod{}_{N}{\mathscr{A}}_{g_{i},n_{i}}$
Let $Y=\tau(X)$. Theorem 5.2 shows that $\chi(Y,K^{\prime})\geq 0$ for any $K^{\prime}\in\operatorname{Perv}(Y)$. This fact together with the claim and theorem 6.1 implies (8.4.1).
Corollary 8.5.
Let ${\mathscr{M}}={\mathscr{M}}_{g_{1},n_{1}}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}$, where either $g_{i}>1$ or $g_{i}=1$ and $n_{i}>0$. If $P$ is perverse sheaf on ${\mathscr{M}}$, then
$\chi^{orb}({\mathscr{M}},P)\geq 0$
Proof.
Let ${\mathscr{M}}^{c}={\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{c}$. The embedding $j:M\to M^{c}$ is affine, since the complement is a divisor. Therefore $j_{*}P\in\operatorname{Perv}({\mathscr{M}}^{c})$. Consequently,
$\chi^{orb}({\mathscr{M}},P)=\chi^{orb}({\mathscr{M}}^{c},j_{*}P)\geq 0$
We note that the previous theorem is false when the $g_{i}=0$. For example, ${\mathscr{M}}_{0,4}^{c}\cong{\mathbb{P}}^{1}$ is a fine moduli space with positive Euler characteristic. Nevertheless, the corollary does hold in this case, and the proof is elementary.
Proposition 8.6.
Let $X={\mathscr{M}}_{0,n_{1}}\times\ldots\times{\mathscr{M}}_{0,n_{r}}$, where all $n_{i}\geq 3$. If $P\in\operatorname{Perv}(X)$, then $\chi(X,P)\geq 0$.
Proof.
Since ${\mathscr{M}}_{0,3}$ is a point, we may assume that $n_{i}>3$. We note that ${\mathscr{M}}_{0,4}\cong{\mathbb{P}}^{1}-\{0,1,\infty\}$. If $E=0+1+\infty$, then $\Omega_{{\mathscr{M}}_{0,4}}^{1}(\log E)$ is nef. Therefore the proposition holds for $X=({\mathscr{M}}_{0,4})^{r}$ by lemmas 4.4 and 4.5. We have an open immersion $j:{\mathscr{M}}_{0,n}\to({\mathscr{M}}_{0,4})^{n-3}$ sending
$(C,p_{1},\ldots,p_{n})\mapsto((C,p_{1},p_{2},p_{3},p_{i}))_{i=4,\ldots n}$
The complement of $\operatorname{Image}(j)$ is given by a (big) diagonal, which is a divisor. Therefore $j$ is affine. It follows that the product map
[…]
is also an open affine embedding. Therefore $J_{*}P$ is perverse when $P\in\operatorname{Perv}(X)$. Therefore $\chi(X,P)\geq 0$. ∎
Remark 8.7.
We note that Theorem 1.2 stated in the introduction follows from Theorem 5.3, Corollary 8.4 and Proposition 8.6.
9. Remarks in characteristic $p$.
In this section, we discuss some pecularities in the characteristic $p$ setting. In particular, we show that the analog of Theorem 5.2 is false in this setting. We fix $N\geq 3$, and consider the moduli space $\prescript{}{N}{{\mathscr{A}}}_{g}$, viewed as a scheme over $\operatorname{Spec}{\mathbb{Z}}[1/N]$. Recall, that this is a smooth scheme over the base. Let $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ (resp. $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}}$) denote the corresponding moduli space over ${\mathbb{F}}_{p}$ (resp. ${\mathbb{C}}$). For a fixed prime $p$, let $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p})$ denote the usual $\ell$-adic Euler characteristic (over $\bar{{\mathbb{F}}}_{p}$) for some fixed prime $\ell\neq p$. Note that this is and integer independent of $\ell$. Similarly, let $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}})$ denote the usual topological Euler characteristic (say with rational coefficients). Note that this can be identified with the $\ell$-adic Euler charactersitic by the usual comparison theorem for étale cohomology and singular cohomology. We begin by first observing the following:
Theorem 9.1.
For all $p\nmid N$, $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p})=\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}}).$
Proof.
Choose a good compactification (over ${\mathbb{Z}}[1/N]$) (for example, a toroidal compactification as in [10]) and use the lemma below. ∎
The following Lemma is standard, but we give a proof for lack of a reference.
Lemma 9.2.
Let $S$ be a strictly henselian ring with with closed point $s$ and generic point $\eta$. Let $\bar{X}\rightarrow S$ a smooth proper morphism, $D\subset\bar{X}$ a relative normal crossings divisor, and $X:=X\setminus D$. Then $\chi(X_{s})=\chi(X_{\bar{\eta}})$.
Proof.
Note that the analogous assertion for $\bar{X}$ holds as a corollary to the classical smooth and proper base change theorem. Suppose $D=D_{1}\cup\cdots\cup D_{n}$ is a decomposition into irreducible component, and let $D_{I}:=\bigcap_{i\in I}D_{i}$ for $I\subset\{1,\ldots,n\}$. Then each $D_{I}\rightarrow S$ is smooth proper, and therefore one also has $\chi((D_{I})_{s})=\chi((D_{I})_{\bar{\eta}})$. By descent, we conclude that $\chi(D_{s})=\chi(D_{\bar{\eta}})$. Here we use the stronger property that the Betti numbers are the same. Finally, by additivity of Euler characteristics we get the desired conclusion for $X$. ∎
Remark 9.3.
We note that for any smooth subvariety of $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ whose closure has normal crossings with $D$ (so that in particular the closure is smooth) and lifts to characteristic 0, we obtain the signed Euler char property as a consequence of the previous Lemma.
We now recall some counter-examples in characteristic p resulting from super-singular strata. Below, for a fixed prime $p$, we set $\prescript{}{N}{{\mathscr{A}}}_{g,p}:=\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}[\zeta_{N}]$ (where $\zeta_{N}$ is a fixed primitive $N$-th root of unity). Note that if $N=1$, we will denote this simply by ${\mathscr{A}}_{g,p}$. We denote by $|\prescript{}{N}{{\mathscr{A}}}_{g,p}|$ the corresponding coarse modulii spaces. If $N>2$, then this is the same as the fine moduli space.
Let $\prescript{}{N}{{\mathcal{S}}}_{g,p}\subset|\prescript{}{N}{{\mathscr{A}}}_{g,p}|$ denote the super singular locus. By definition this is the locus of points $(A,\lambda,H)$ (here $\lambda$ is a principal polarization and $H$ is a level structure) such that $A\otimes\bar{{\mathbb{F}}}_{p}$ is supersingular. This is a closed subset with $\prescript{}{N}{\dim}({\mathcal{S}}_{g,p})=\lfloor\frac{g^{2}}{4}\rfloor$ ([22] ). Note that it is known that in general $|{\mathcal{S}}_{g,p}|$ is not reducible, but it is equidimensional. In particular, $\dim(\prescript{}{N}{{\mathcal{S}}}_{2,p})=1$ and $\dim({\mathcal{S}}_{3,p})=2$. The components of ${\mathcal{S}}_{g,p}$ can be parametrized by spaces flag type data for Dieudonne modules ([22]). We do not recall the details, but give a description of the resulting parameter space in the case $g=2$ ([17] (section 2, 5.1, 5.3) and $g=3$ [22]
Theorem 9.4.
([17],[22])
  1. (1)
    The irreducible components of $|\prescript{}{N}{{\mathcal{S}}}_{2,p}|$ are isomorphic to ${\mathbb{P}}^{1}$.
  2. (2)
    Let ${\mathcal{X}}\subset{\mathbb{P}}^{2}_{{\mathbb{F}}_{p}}$ denote the Fermat curve of degree $p+1$ (i.e. defined by $x^{p+1}+y^{p+1}+z^{p+1}=0$), and ${\mathcal{Y}}:={\mathbb{P}}({\mathcal{O}}_{{\mathcal{X}}}(1)\oplus{\mathcal{O}}_{{\mathcal{X}}}(-1))$. There is a natural morphism
    $\pi:{\mathcal{Y}}\rightarrow{\mathcal{S}}_{3,p}$
    such that:
    1. (a)
      $\pi$ is of degree one onto its image.
    2. (b)
      The image of $\pi$ is an irreducible component of ${\mathcal{S}}_{3,p}$.
    3. (c)
      Let ${\mathcal{D}}\subset{\mathcal{Y}}$ denote the the unique section with negative self intersection. Then the morphism $\pi$ contracts ${\mathcal{D}}$ to a point.
Theorem 9.5.
  1. (1)
    There is a smooth projective curve $C\subset\prescript{}{N}{{\mathscr{A}}}_{2,p}$ which does not satisfy the signed Euler characteristic property i.e. the $\ell$-adic Euler charactersitic $\chi(C_{\bar{{\mathbb{F}}}_{p}})>0$. In fact, $C={\mathbb{P}}^{1}$.
  2. (2)
    Suppose $p>2$. There exists a normal surface in ${\mathscr{A}}_{3,p}$ with isolated singularities which does not satisfy the signed Euler characteristic property.
Proof.
The first part is a direct consequence of the first part of the previous theorem. For the second part, we note that the blow down of the surface given in part two satisfies the properties stated here. Namely, it is a normal surface with isolated singular point, and it has negative Euler characteristic. This follows from the fact that the Fermat curve in this case has negative Euler characteristic, and therefore so does the projective bundle described above. ∎
Corollary 9.6.
  1. (1)
    There is an embedding $({\mathbb{P}}^{1})^{[g/2]}\subset\prescript{}{N}{{\mathscr{A}}}_{g,p}$ for all $g\geq 2$.
  2. (2)
    There exists a sequence of smooth subvarieties $V_{g_{i}}\subset{}_{N}A_{g_{i},p}$, such that $(-1)^{\dim V_{g_{i}}}\chi(V_{g_{i}})\to-\infty$.
  3. (3)
    The contangent bundle $\Omega^{1}{\prescript{}{N}{{\mathscr{A}}}_{g,p}}$ and the Hodge bundle $F^{1}$ associated to the universal family of abelian varieties on $\prescript{}{N}{{\mathscr{A}}}_{g,p}$ are not nef (in the sense that their restrictions to suitable projective subvarieties are not nef).
Proof.
If $g=2$, (1) is the first part of the previous theorem. Suppose $g>2$, and let $d=[g/2]$. If $g$ is odd, let $E$ be a fixed polarized elliptic curve with level $N$ structure. Then an irreducible component of either
$\{A_{1}\times\ldots\times A_{d}|A_{i}\in{}_{N}{\mathcal{S}}_{2,p}\}\quad\text{if $g$ even}$
$\{A_{1}\times\ldots\times A_{d}\times E|A_{i}\in{}_{N}{\mathcal{S}}_{2,p}\}\quad\text{if $g$ even}$
gives an embedded copy of $({\mathbb{P}}^{1})^{d}$ in ${}_{N}{\mathscr{A}}_{g,p}$.
(2) follows immediately from (1).
For the last item, observe that the restriction of $\Omega^{1}_{{}_{N}{\mathscr{A}}_{g,p}}$ to a smooth rational projective curve is not nef, because it has $\mathcal{O}(-2)$ as a quotient. One has that
$\Omega^{1}_{{}_{N}{\mathscr{A}}_{g,p}}\cong S^{2}F^{1}$
e.g. by [10, chap IV, theorem 7.7]. Therefore $F^{1}$ cannot be nef.
Remark 9.7.
  1. (1)
    The last corollary shows that Theorem 1.2 fails very badly in positive characteristic.
  2. (2)
    The corollary gives a counterexample to Theorem 2.2 on the semipositivity (i.e. nefness) of Hodge bundles in characteristic p. This goes back to Moret-Bailly [25] who used essentially the same example when $g=2$.